OpenStudy (elisaneedshelp):

Can some1 help me find the apothem and face of a regular dodecahedron? Please por favor

OpenStudy (elisaneedshelp):

@mathslover

OpenStudy (elisaneedshelp):

@ganeshie8 plz

ganeshie8 (ganeshie8):

are u given the side of polygon ?

ganeshie8 (ganeshie8):

or any other info given ?

OpenStudy (elisaneedshelp):

the edge is 3.3 and the height of Int pyramid is 3.75 (for volume)

ganeshie8 (ganeshie8):

hey could u post the snapshot of original quesiton if psble

OpenStudy (elisaneedshelp):

yea sure, sorry for the delay my mom was calling me to do something and i had no choice

ganeshie8 (ganeshie8):

lol okay :) i think we're gona mess wid pentagon

ganeshie8 (ganeshie8):

http://en.wikipedia.org/wiki/Dodecahedron

OpenStudy (elisaneedshelp):

actually now that i knocked some sense into my brain, there is no actual problem lol thats all that i was given and i was able to find the surface area and volume but not the apothem or the face :/

ganeshie8 (ganeshie8):

if you have surface area, just divide 12 to get the area of face ?

OpenStudy (elisaneedshelp):

so then it would be 18.73599889

OpenStudy (elisaneedshelp):

18.74

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

did u use some formula for surface area ?

OpenStudy (elisaneedshelp):

yeah i used SA= 3 sqrt |25+10|sqrt5| a^2

ganeshie8 (ganeshie8):

good, now that u have area of face, can u find apothem ?

OpenStudy (elisaneedshelp):

i really dont understand how to get apothem :(

ganeshie8 (ganeshie8):

use below : Area of face = 1/2 * (perimeter) * (apothem)

ganeshie8 (ganeshie8):

since the face is a pentagon, it has 5 edges. so perimeter = ?

OpenStudy (elisaneedshelp):

16.5 since each edge is 3.3

ganeshie8 (ganeshie8):

yes, plug that

ganeshie8 (ganeshie8):

Area of face = 1/2 * (perimeter) * (apothem) 18.74 = 1/2 * 16.5 * (apothem)

ganeshie8 (ganeshie8):

you can solve `apothem`

OpenStudy (elisaneedshelp):

would it be 2.271515152?

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (elisaneedshelp):

would the volume for this shape be 275.39 ?

ganeshie8 (ganeshie8):

volume = 275.39 is correct

OpenStudy (elisaneedshelp):

is it possible for you to help me with three more shapes and then i wont bother you anymore :) and thanks

ganeshie8 (ganeshie8):

sure, il try :)

OpenStudy (elisaneedshelp):

yay thanks can you help me find the volume of a triangular prism with the info given: ~Length of rectangle:7 ~base of triangle:6.4 ~base of angle triangle:63 degrees

ganeshie8 (ganeshie8):